OpenStudy (anonymous):

Find the inverse of y=2√(x+3). Thanks!

OpenStudy (nubeer):

y = f(x) square both sides. y^2=4(x+3). now u have to make the x alone y^2 / 4 =x+3 (y^2 / 4 ) - 3 = x \[f^{-1}(x)\] = (x^2/4) - 3

OpenStudy (anonymous):

thanks! I wasn't sure how to get rid of the square root. :)

OpenStudy (nubeer):

no problem at all. :) just square whenever u vanish the square root.